OpenStudy (anonymous):

Daniel collected 120 sports cards. Sixty-five of the cards are baseball cards. About what percent of the cards are baseball cards? Round to the nearest percent, if necessary. A. 65% B. 54% C. 40% D. 46%

OpenStudy (anonymous):

120 - 65 = 55 55 / 120 = 0.46 So 46% of the cards are baseball cards.

OpenStudy (anonymous):

To give a medal click "Best Response".

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

(65/120)*100

OpenStudy (anonymous):

what is the proportion? 9/x = 4/12 A. 17 B. 13 C. 27 D. 3

OpenStudy (anonymous):

54%

OpenStudy (anonymous):

9/x=4/12 9/x=1/3 x=27

OpenStudy (anonymous):

The rectangles shown are similar. What is the value of x? A. 4 B. 6 C. 5 D. 9

OpenStudy (anonymous):

one